Determine if the outcome is unusual. Consider as unusual any result that differs from the mean by more than 2 standard deviations. That is, unusual values are either less than μ - 2σ or greater than μ + 2σ.

The Acme Candy Company claims that 60% of the jawbreakers it produces weigh more than .4 ounces. Suppose that 800 jawbreakers are selected at random from the production lines. Would it be unusual for this sample of 800 to contain 423 jawbreakers that weigh more than .4 ounces?

Yes

No
Answer #1

Solution

given : p=0.60, n = 800

Mean (u) = np= 800 * 0.60 = 480

Standard deviation (0) = V np(1-P) = V800 * 0.60 * (1 – 0.60) = 13.86

u – 20 = 480 – (2 * 13.86) = 452.28

u + 20 = 480 + (2 * 13.86) = 507.72

Unusual values are either less than u – 20 or greater than u + 20

Yes, since 423 is less than u – 20 (that is 452.28)